# TEncDM: Understanding the Properties of Diffusion Model in the Space of Language Model Encodings

### v1

@@ 6 Datasets
To evaluate the performance of our diffusion models we use three datasets in English language. The ROCStories Mostafazadeh et al. (2016) dataset contains 98k five-sentence commonsense fictional stories, that capture causal and temporal relations between daily events. The subset of QQP Chen et al. (2017) dataset, proposed in Gong et al. (2023), consists of 144k question pairs from the Quora platform that are paraphrases of each other. The XSum Narayan et al. (2018) dataset is used for summarization problem and it contains 204k BBC articles, which are provided as document and summary pairs [^1] All the datasets we use in this work are publicly available under a creative commons or an open source license.. The detailed statistics for each dataset can be found in Appendix F.
@@ 7 Empirical Analysis
In this section, we evaluate the components of our framework on the ROCStories dataset. To simplify the setup, we only consider unconditional generation. In Section 8, we demonstrate that our findings can be successfully transferred to the conditional generation problems. In this section, we do not compare our method with others. The comparison with the GPT2 is presented in Appendix G.
@@ 7.1 Evaluation Metrics
We follow the model evaluation scheme from the (Lovelace et al., 2022). To evaluate the quality of our model we use Perplexity (ppl), calculated with GPT-2 Large Radford et al. (2019). To measure the diversity of the generated text we utilize the diversity metric proposed in Su et al. (2022). We calculate it as div$(y)=\prod_{n=2}^{4}\frac{|\text{\# of unique n-grams in }y|}{|\text{\# of n-grams in }y|}$, where $y$ is a set of generated texts. To ensure that the model does not reproduce the training dataset during the generation we evaluate the Memorization (mem). We calculate it as the proportion of generated 4-grams that are found in the training set. As Perplexity tends to be small for the texts with repetitions, we also use MAUVE Score Pillutla et al. (2021) to estimate the quality of text. MAUVE is a language model-based metric that measures the distance between the distributions of generated and reference texts using divergence frontiers. We leave all MAUVE hyperparameters at the default values presented in the original paper.
@@ 7.2 Model setup
The training of our diffusion model is conducted within the latent space of BERT encodings, as it has shown the best performance among all encoders. We employ a 3-layer transformer for the decoder and train it to reconstruct $z_{0}$ from $z_{t}$, where $t\in U[0,0.15]$. A comprehensive analysis of various decoder modifications is presented in Section 7.4 and Appendix B. The diffusion model is the 12-layer transformer with dimensionality of 768. By default we train it with tan-9 noise scheduler.
Encoder ppl $\downarrow$ mem $\downarrow$ div $\uparrow$ mauve $\uparrow$
BERT emb $48.9_{.36}$$371_{.003}.$$324_{.002}.$$600_{.016}.$
BERT $\textbf{34.1}_{.66}$$412_{.005}.$$304_{.006}.$$\textbf{.707}_{.024}.$
T5 $47.7_{.66}$$361_{.001}.$$330_{.001}.$$475_{.008}.$
Table 1: Comparison of diffusion encoders.
7.3 Effect of Diffusion Encoder
@@
We compare latent spaces of BERT (Devlin et al., 2019) (bert-base-cased) and T5 (Raffel et al., 2020) (t5-base) encodings, as well as BERT embeddings, to ascertain the optimal choice for the diffusion model. In this experiment, we train diffusion models with the same set of hyperparameters across all diffusion encoders. We train the decoders according to the scheme described in Section 7.2. The results of this comparison are presented in Table 1 and they show a clear advantage of the latent space derived from BERT encodings. div and mem for T5 encoder and BERT embeddings are better, because their generated texts include words that do not fit the context. The text samples are presented Table 9 of Appendix H. This confirms our hypothesis that encodings are better suited for the training of a diffusion model.
7.4 Effect of Decoder
@@
To confirm the hypothesis about the importance of the decoder architecture and its training scheme, we compare an MLP decoder consisting of two linear layers with a 3-layer transformer. We corrupt the decoder input $z_{0}$ by transforming it into $z_{t}$, using the diffusion forward process with $t\in U[0,0.15]$. We choose this method, because it brings the decoder input closer to the diffusion output. A more detailed analysis of corruption techniques is presented in the Appendix B. To keep the experiment fair, we apply all decoders to the same generated latents. The results of the experiment are shown in Table 2. The MLP decoder achieves the worst text quality, because it overfits on the special token embeddings and fails to decode them from the generated latents. Examples of the generated samples are shown in Appendix H. Corruption of the input helps to avoid overfitting. At the same time, incorporating contextual information into the decoder increases the quality even more
Decoder ppl $\downarrow$ mem $\downarrow$ div $\uparrow$ mauve $\uparrow$
MLP $607.1_{15.6}$$332_{.003}.$$400_{.004}.$$004_{.00}.$
  + $Cor(z_{0})$ $36.2_{1.8}$$415_{.005}.$$301_{.006}.$$650_{.03}.$
Transformer $40.4_{.86}$$408_{.005}.$$308_{.006}.$$568_{.02}.$
  + $Cor(z_{0})$ $\textbf{34.1}_{.66}$$412_{.005}.$$304_{.006}.$$\textbf{.707}_{.02}.$
Table 2: Comparison of decoders for encoding-based diffusion model.
7.5 Effect of self-conditioning
We conduct a series of experiments to understand how self-conditioning affects the diffusion model. In Figure 2, we compare the quality of the models with and without self-conditioning for different number of denoising steps. The results show that while the quality of the model without self-conditioning increases as the number of steps increases, the quality of the model with self-conditioning reaches a maximum at a value of 50 steps in terms of MAUVE, after which it starts to drop. Nevertheless, at the highest point model with self-conditioning surpasses the model without it according to both MAUVE and perplexity.
We explain this drop in generation quality with mismatch between diffusion model inputs at train and inference stages. To confirm our hypothesis, we calculated the mean-squared norm (magnitude) of the values of each latent $\hat{z}_{0}^{t}$ in a mini-batch predicted by the diffusion model during generation (i.e. $\frac{1}{N\cdot d\cdot m}\|\hat{z}_{0}^{t}\|_{2}^{2}$, where $N$ is a batch size, $d$ is a dimension and $m$ is a sequence length). We plot this magnitude with respect to timestep for generations with different number of steps as well as for the predictions $\bar{z}_{0}^{t}$ from the training stage. The results are presented in Figure 3. They indicate that self-conditioning significantly increases the prediction magnitude as the number of steps increases. This can be explained by the following: during training, the model learns to use self-conditioning to approximate $z_{0}$ more accurately. Consequently, self-conditioning increases the model’s confidence, which is directly related to prediction magnitude. During the generation process, the model takes its own prediction, which has an increased magnitude, as an input at each step and increases it further. Therefore, the increase in magnitude depends directly on the number of generation steps. Eventually, this leads to a mismatch between the predictions fed into the model during training and generation. In the Appendix C, we provide a more detailed discussion of this phenomenon. It is worth noting that the smallest mismatch is observed for the trajectory of 50 generation steps, which corresponds to the best quality.
@@ 7.5 Effect of self-conditioning
Figure 2: Comparison of models with and without self-conditioning
@@
Figure 3: Comparison of magnitudes for generation processes with different amount of steps.
7.6 Effect of Noise scheduler
@@
While the sqrt noise scheduler adds significantly more amount of noise in the initial timesteps than cosine one, the rate of noise addition decreases for the subsequent timesteps. As a result, the denoising task becomes insufficiently hard for the timesteps $t\in[0,0.5]$, which should lead to a decrease in their contribution to the generation process. This can be seen from the reconstruction accuracy. In contrast, tan-d noise scheduler adds more noise consistently across all timesteps, leading to a more challenging training task and improved generation performance.
Based on these observations, we conclude that in order to improve the efficiently of the denoising process, it is essential to increase the amount of added noise within all timesteps. However, it is important to strike a balance as adding excessive noise can negatively impact performance. In our experiments, tan-9 produces the best result in terms of mauve keeping the mem and div reasonable.
As a rule of thumb, the noise schedule should be such that the diffusion model recovers approximately the same amount of information at each timestep. Otherwise, some of the them will not contribute to the denoising process enough.
Figure 4: Comparison of noise schedulers.
@@
tan-7 $34.4_{.77}$$395_{.004}.$$320_{.002}.$$688_{.023}.$
tan-9 $34.1_{.66}$$412_{.005}.$$304_{.006}.$$\textbf{.707}_{.024}.$
tan-11 $31.9_{.31}$$428_{.004}.$$288_{.003}.$$694_{.026}.$
tan-13 $35.5_{.62}$$406_{.003}.$$298_{.002}.$$676_{.031}.$
@@ 8 Seq2Seq Experiments
Method Sampling R-L $\uparrow$ BS $\uparrow$ B-4 $\uparrow$
${}^{\dagger}$ Random $52.7$ $82.4$
${}^{\dagger}$ $82.9$ $23.3$
${}^{\dagger}$ (BERT) 56.4 83.0 30.4
${}^{\dagger}$ (T5) 52.4 81.6 26.4
${}^{\dagger}$ MBR-10 58.8 $83.7$ $24.1$
${}^{\dagger}$ 84.0 $24.3$
TEncDM (${}^{\dagger}$ $58.1$ 84.0 31.8
TEncDM (${}^{\dagger}$ $53.5$ $82.3$ $27.4$
GPT2-small ${}^{\star}$ Nucleus $52.1$ $82.5$ $19.8$
${}^{\star}$ $57.5$ $83.8$ $27.2$
Table 4: Seq2Seq evaluation results of AR and Diffusion methods on QQP. We calculate ROUGE-L (R-L), BERTScore (BS) and BLEU-4 (B-4).
Method Sampling ROUGE-1/2/L $\uparrow$ BS $\uparrow$
${}^{\diamond}$ 24.0 / 3.9 / 20.3
${}^{\diamond}$ 24.0 / 4.0 / 20.3
${}^{\diamond}$ 23.8 / 3.6 / 20.2
${}^{\diamond}$ 24.8 / 4.2 / 20.8
${}^{\dagger}$ Random 18.9 / 1.3 / 13.6 $46.8$
TEncDM (${}^{\dagger}$ Random 32.2 / 10.8 / 25.7 69.5
TEncDM (${}^{\dagger}$ Random 32.4 / 10.9 / 25.7 68.8
${}^{\dagger}$ MBR-5 $19.3$ / $1.7$ / $14.1$ $46.9$
TEncDM (${}^{\dagger}$ MBR-5 32.8 / 11.2 / 26.2 69.8
TEncDM (${}^{\dagger}$ MBR-5 32.9 / 11.4 / 26.5 69.2
${}^{\dagger}$ MBR-50 $29.3$ / $8.3$ / $24.7$
${}^{\dagger}$ MBR-50 $31.7$ / $10.1$ / $24.7$
${}^{\star}$ Nucleus $30.5$ / $10.4$ / $24.2$
Table 5: Seq2Seq evaluation results of NAR, AR and Diffusion methods on XSum. BS is a BERTScore.
We conduct experiments to validate the effectiveness of the proposed method on two different tasks, against ten AR ($\star$), non-diffusion NAR ($\circ$) and diffusion NAR ($\dagger$) baselines.
@@
For evaluation of paraphrasing task, we adopt the setting of SeqDiffuSeq Yuan et al. (2022) and calculate ROUGE-L Lin (2004), BERTScore Zhang et al. (2019) and BLEU-4. In addition, we follow the approach of Wu et al. (2023) and report ROUGE-1/2 for summarization task.
Baselines
We include three groups of baselines. The first group comprises of classical AR baselines: Transformer Vaswani et al. (2017b) and finetuned GPT-2 Radford et al. (2019). We also compare against NAR methods: NAT Gu et al. (2017), iNAT Lee et al. (2018), CMLM Ghazvininejad et al. (2019), LevT Gu et al. (2019). Besides, we compare the approach to other diffusion-based methods: DiffuSeq Gong et al. (2023), SeqDiffuSeq Yuan et al. (2022), GENIE Lin et al. (2023), AR-diffusion Wu et al. (2023).
@@ Results
We report our comparison on QQP and XSum in Table 4 and Table 5, respectively. We took the results of NAR and AR approaches from the corresponding papers Qi et al. (2021); Wu et al. (2023); Yuan et al. (2022).
We use BERT as diffusion encoder and experiment with two conditional encoders: BERT and T5. We observe that both encoders are effective for XSum and QQP datasets, but using BERT leads to a better quality on QQP across all metrics and on XSum these encoders performs similarly.
The comparison with other methods clearly demonstrate that TEncDM outperforms the existing non-diffusion NAR approaches across all metrics. Furthermore, TEncDM surpasses diffusion and AR approaches by a large margin on summarization task. It also achieves consistent improvements over diffusion models on QQP with random candidate sampling.
Recent works Li et al. (2022); Wu et al. (2023) utilize Minimum Bayes Risk (MBR) Kumar and Byrne (2004) decoding to select the best sample. For fair comparison, we also employ MBR decoding with the same number of candidates. As we can see from Table 5, TEncDM significantly outperforms diffusion baselines with even less number of candidates on XSum. At the same time, Table 4 shows that the results on QQP are comparable with other models.

### v2

@@ 6 Datasets
To evaluate the performance of our diffusion models we use five datasets in English language. Two of them are unconditional: ROCStories and Wikipedia, and three are conditional: QQP, XSum and Wiki-Auto. The ROCStories (Mostafazadeh et al. 2016) dataset contains 98k five-sentence commonsense fictional stories, that capture causal and temporal relations between daily events. The Wikipedia dataset is obtained from the ROOTS corpus (Laurençon et al. 2023), it is a collection of over 2 million cleaned articles from the Wikipedia platform. The subset of QQP (Chen et al. 2017) dataset, proposed in (Gong et al. 2023), consists of 144k question pairs from the Quora platform that are paraphrases of each other. The XSum (Narayan, Cohen, and Lapata 2018) dataset is used for summarization problem and it contains 204k BBC articles, which are provided as document and summary pairs. The Wiki-Auto (Jiang et al. 2020) dataset consists of aligned sentences from complex and simplified Wikipedia [^1] All the datasets we use in this work are publicly available under a creative commons or an open source license.. The detailed statistics for each dataset can be found in Appendix F.
@@ 7 Empirical Analysis
Decoder ppl $\downarrow$ mem $\downarrow$ div $\uparrow$ mauve $\uparrow$
ROCStories
MLP $39.7_{3.38}$$444_{.002}.$$297_{.004}.$$716_{.074}.$
+ $Cor(z_{0})$ $31.2_{.33}$$448_{.002}.$$293_{.003}.$$739_{.051}.$
Transformer $34.2_{.29}$$445_{.001}.$$295_{.003}.$$714_{.037}.$
+ $Cor(z_{0})$ $\textbf{29.1}_{.89}$${.453}_{.003}.$${.295}_{.002}.$$\textbf{.762}_{.043}.$
Wikipedia
Transformer $180.6_{3.2}$$261_{.001}.$$511_{.001}.$$526_{.025}.$
+ $Cor(z_{0})$ $\textbf{104.4}_{2.1}$$286_{.002}.$$504_{.003}.$$\textbf{.874}_{.011}.$
Table 2: Comparison of decoders for encoding-based diffusion model.
In this section, we evaluate the components of our framework on the ROCStories and Wikipedia datasets. To simplify the setup, we only consider unconditional generation. In Section 8, we demonstrate that our findings can be successfully transferred to the conditional generation problems. In this section, we do not compare our method with others. The comparison with the GPT-2 (Radford et al. 2019) on unconditional generation is presented in Appendix I.
@@ 7.1 Evaluation Metrics
We follow the model evaluation scheme from the (Lovelace et al. 2023). To evaluate the quality of our model we use Perplexity (ppl), calculated with GPT-2 Large (Radford et al. 2019). To measure the diversity of the generated text we utilize the diversity metric proposed in (Su et al. 2022). We calculate it as div$(y)=\prod_{n=2}^{4}\frac{|\text{\# of unique n-grams in }y|}{|\text{\# of n-grams in }y|}$, where $y$ is a set of generated texts. To ensure that the model does not reproduce the training dataset during the generation we evaluate the Memorization (mem). We calculate it as the proportion of generated 4-grams that are found in the training set. As Perplexity tends to be small for the texts with repetitions, we also measure MAUVE Score (Pillutla et al. 2021) to estimate the quality of text. MAUVE is a language model-based metric that measures the distance between the distributions of generated and reference texts using divergence frontiers. We leave all MAUVE hyperparameters at the default values presented in the original paper.
@@ 7.2 Model setup
The training of TEncDM is conducted within the latent space of BERT encodings, as it has shown the best performance among all encoders. We employ a 3-layer transformer for the decoder and train it to reconstruct $z_{0}$ from $z_{t}$, where $t\in U[0,0.15]$. A comprehensive analysis of various decoder modifications is presented in Section 7.2 and Appendix B. The diffusion model is the 12-layer transformer with dimensionality of 768. By default we train it with tan-9 noise scheduler.
Effect of Diffusion Encoder
@@
We compare latent spaces of BERT (Devlin et al. 2019), RoBERTa (Liu et al. 2019), BART (Lewis et al. 2020) and T5 (Raffel et al. 2020) encoders, as well as BERT embeddings, to ascertain the optimal choice for the diffusion model. All encoders have approximately the same size of 100M parameters. In this experiment, we train diffusion models with the same set of hyperparameters across all diffusion encoders. We train the decoders according to the scheme described in Section 7.2. The results of this comparison are presented in Table 1 and they show a clear advantage of the latent space derived from BERT encodings on ROCStories dataset. Furthermore, the quality of all encoders is superior to that of BERT embeddings. A better div and mem for embeddings can be explained by the presence of words in the corpus that do not align with the context. The text samples are presented Table 9 of Appendix J. This confirms our hypothesis that encodings are better suited for the training of a diffusion model. We discuss the drop in mauve for RoBERTa in Appendix G.
Effect of Decoder
@@
To confirm the hypothesis about the importance of the decoder architecture and its training scheme, we compare an MLP decoder consisting of two linear layers with a 3-layer transformer. The latter is able to extract contextual information for each token. We corrupt the decoder input $z_{0}$ by transforming it into $z_{t}$, using the diffusion forward process with $t\in U[0,0.15]$. We choose this method, because it brings the decoder input closer to the diffusion output. A more detailed analysis of corruption techniques is presented in the Appendix B. To keep the experiment fair, we apply all decoders to the same generated latents. The results of the experiment are shown in Table 2. The MLP decoder without corruption achieves the lowest text quality in terms of perplexity, but comparable by mauve with Transformer without corruption. However, it is challenging to make meaningful comparisons between decoders by mauve due to the significant variance. From this experiment, we can conclude that corruption of the latent helps to improve the quality for both datasets. At the same time, the incorporation of contextual information into the decoder lead to the best result.
@@
Figure 2: Comparison of models with and without self-conditioning on ROCStories dataset.
@@
Figure 3: Comparison of prediction magnitudes for generation processes with different amount of steps on ROCStories dataset.
Effect of self-conditioning
We conduct a series of experiments to understand how self-conditioning (SC) affects the denoising process. In Figure 2, we compare the quality of the models with and without SC for different number of denoising steps on the ROCStories dataset. The results show that while the quality of the model without SC increases as the number of steps increases, the quality of the model with SC reaches a maximum at a value of 50 steps in terms of mauve, and then it starts to drop. Nevertheless, at the highest point the model with SC surpasses the model without it according to both mauve and perplexity.
We explain this drop in generation quality with mismatch between diffusion model inputs at train and inference stages. To confirm our hypothesis, we calculated the mean-squared norm (magnitude) of the values of each latent $\hat{z}_{0}^{t}$ in a mini-batch predicted by the diffusion model during generation (i.e. $\frac{1}{N\cdot d\cdot m}\|\hat{z}_{0}^{t}\|_{2}^{2}$, where $N$ is a batch size, $d$ is a dimension and $m$ is a sequence length). We plot this magnitude with respect to timestep for generations with different number of steps as well as for the predictions $\bar{z}_{0}^{t}$ from the training stage. The results for the ROCStories dataset are presented in Figure 3. They indicate that self-conditioning significantly increases the prediction magnitude as the number of steps increases. This can be explained by the following: during training, the model learns to use self-conditioning to approximate $z_{0}$ more accurately. Consequently, self-conditioning increases the model’s confidence, which is directly related to prediction magnitude. During the generation process, the model takes its own prediction, which has an increased magnitude, as an input at each step and increases it further. Therefore, the increase in magnitude depends directly on the number of generation steps. Eventually, this leads to a mismatch between the predictions fed into the model during training and generation.
In the Appendix C, we provide a more detailed discussion of this phenomenon and show that the same behavior is observed in the larger Wikipedia dataset. It is worth noting that the smallest mismatch is observed for the trajectory of 50 generation steps, which corresponds to the best quality.
Figure 4: Comparison of noise schedulers on ROCStories dataset.
Effect of Noise scheduler
@@
While the sqrt noise scheduler adds significantly larger amount of noise in the initial timesteps than cosine one, the rate of noise addition decreases for the subsequent timesteps. For both schedulers, the denoising task becomes insufficiently hard for the most timesteps, which leads to a decrease in their contribution to the generation process. This can be seen from the reconstruction accuracy. In contrast, tan-d noise scheduler adds more noise consistently across all timesteps, leading to a more challenging training task and improved generation performance.
Based on these observations, we conclude that in order to improve the efficiently of the denoising process, it is essential to increase the amount of added noise within all timesteps. However, it is important to strike a balance as adding excessive noise can negatively impact performance. In our experiments, tan-9 yielded results that were marginally superior to all other schedulers in terms of mauve, with slightly better memorization and diversity, while lagging behind tan-7 in terms of perplexity.
As a rule of thumb, the noise schedule should be such that the diffusion model recovers approximately the same amount of information at each timestep. Otherwise, some of them will not contribute to the denoising process enough.
@@
tan-7 $\textbf{27.1}_{.31}$$455_{.004}.$$286_{.001}.$$730_{.026}.$
tan-9 $29.1_{.89}$${\textbf{.453}}_{.003}.$${\textbf{.295}}_{.002}.$$\textbf{.762}_{.043}.$
tan-11 $33.4_{2.4}$$464_{.006}.$$279_{.004}.$$730_{.038}.$
Wikipedia
sqrt $364.0_{6.5}$$139_{.001}.$$664_{.004}.$$325_{.037}.$
tan-9 $\textbf{104.4}_{2.1}$$286_{.002}.$$504_{.003}.$$\textbf{.874}_{.011}.$
@@
Method QQP XSum Wiki-Auto
R-L $\uparrow$ BS $\uparrow$ B-4 $\uparrow$ R-1/2/L $\uparrow$ BS $\uparrow$ R-L $\uparrow$ BS $\uparrow$ B-4 $\uparrow$
DiffuSeq 52.7 82.4 18.9 / 1.3 / 13.6 46.8 79.1 26.1
SeqDiffuSeq⋆† 82.9 23.3 14.1 / 1.1 / 11.4 58.4 82.1 37.1
Genie 29.3 / 8.3 / 24.7
AR-Diffusion 54.9 81.4 31.2 27.1 / 6.4 / 20.8 59.7 54.9 81.3 32.7
TEncDM (BERT) 56.4 82.4 30.2 31.5 / 10.0 / 24.9 68.2 58.1 80.5 41.6
TEncDM (T5) 57.3 83.8 30.7 33.4 / 11.4 / 26.8 70.1 57.7 81.2 41.6
TEncDM (RoBERTa) 55.8 82.4 30.0 33.7 / 11.9 / 27.1 69.8 57.9 81.0 40.5
GPT2-small FT $52.1$ $82.5$ $19.8$ $54.6$ $80.2$ $30.8$
Transformer-base $57.5$ $83.8$ $27.2$ $30.5$ / $10.4$ / $24.2$
FLAN-T5-base $52.3$ $83.2$ 34.6 / 12.9 / 27.2 72.7
Table 4: Seq2Seq evaluation results of AR and Diffusion methods on QQP, XSum and Wiki-Auto datasets. We calculate ROUGE-1/2/L (R-1/2/L), BERTScore (BS) and BLEU-4 (B-4). Result of AR approaches, DiffuSeq, and SeqDiffuSeq were sources from their respective publications (marked as $\star$). Additionally, we trained AR-Diffusion and SeqDiffuSeq on previously unreported datasets, adhering to the methodologies outlined by the authors (marked as $\dagger$).
@@ 8 Seq2Seq Experiments
We are conducting experiments to evaluate the effectiveness of the proposed components for text diffusion generation on language model encodings on three different tasks: paraphrasing (QQP), summarization (XSum), and text simplification (Wiki-Auto).
Baselines
We include two groups of baselines in comparison. The first group comprises of classical AR baselines: Transformer (Vaswani et al. 2017a), FLAN-T5-base (Chung et al. 2024) and finetuned GPT-2-small (Radford et al. 2019). Besides, we compare TEncDM to popular embedding-based diffusion methods: DiffuSeq (Gong et al. 2023), SeqDiffuSeq (Yuan et al. 2024), GENIE (Lin et al. 2023), AR-diffusion (Wu et al. 2023). We focus only on non-autoregressive diffusion models trained in the latent space of embeddings or encodings.
@@
For evaluation of paraphrasing and simplification tasks, we adopt the setting of SeqDiffuSeq (Yuan et al. 2024) and calculate ROUGE-L (Lin 2004), BERTScore (Zhang et al. 2020) and BLEU-4. In addition, we follow the approach of Wu et al. (2023) and report ROUGE-1/2/L for summarization task.
@@ Results
Table 4 presents a comprehensive comparison of our approach against existing methods across all datasets. Results for AR approaches, DiffuSeq, and SeqDiffuSeq were sourced from their respective publications (Qi et al. 2021; Wu et al. 2023; Yuan et al. 2024). Furthermore, we trained both AR-Diffusion and SeqDiffuSeq on previously unreported datasets, adhering to the methodologies outlined by the authors.
We experiment with three encoders: BERT-base, T5-base, RoBERTa-base to investigate their efficacy on conditional tasks. We use the same encoder for $E_{diff}$ and $E_{cond}$. Our findings indicate that all three encoders demonstrate effectiveness across the tasks, achieving comparable performance levels. However, no single encoder consistently outperforms the others across all tasks. T5-base excels in question-question paraphrasing (QQP) across all metrics, while RoBERTa-base demonstrates superior performance on extreme summarization (XSum). BERT-base and T5-base exhibit comparable performance on text simplification.
A comparison against other methods clearly demonstrates that TEncDM, utilizing any of the tested encoder models, outperforms popular diffusion embedding-based approaches. Furthermore, TEncDM-T5 achieves either superior or equivalent performance to all classical AR baselines on QQP. Similarly, TEncDM-RoBERTa demonstrates comparable quality against FLAN-T5-base on extreme summarization (XSum). Furthermore, TEncDM surpasses diffusion and AR approaches by a large margin on simplification task.